OpenStudy (anonymous):

hopefully someone can help me out... Perform the operation 6+SR (-18) all divided by 3

OpenStudy (anonymous):

\[(6 +\sqrt{(-18)}) \div 3\]

OpenStudy (radar):

\[(6+\sqrt{9 (-2})\div3\]

OpenStudy (radar):

\[(6+3\sqrt{-2})3\]